1 Suppose that two identical firms produce widgets and that they are the only firms in the market. Their costs are given by CFind the Cournot-Nash equilibrium. Calculate the profit of each firm at this equilibrium. (For all of the following, enter aFirm 1s would earn profit of $ 518,400.00. Returning to the duopoly above, suppose Firm 1 abides by the agreement but Firm 2The answers I filled are wrong.
Answer #1

Cost functions c, = 60 Q. , C2 = 60 Q2 P = 900-9 = 900- (alt &2) Cougnot duopoly. Profit of Firm-1 (Th) - The Tei Pai - 609 (Solving BRI and Bri Simultaneously 9 = 8uo - Q2 2 201 - 8uo- Q2 - 8uo- 840- QI ( 940,00) from BR2 =) 201 2 (840) - suo tal 2each firm will earn Profit of $ 78400 Cartel Profit function of cartel from π= TR - TO Pico - 6o ♡ (P-60 ) 11 (900-9-60 19 TQ = 420 P= 900-8 900 - 420 =480 T = T2 = (P-60) Qi (480-60) (210) (420)(210) 88200 So each from will earn Profit of $ 88 200.duolory with Cheating from-1 abides The agreement by Producing 210 units How much from-2 will produce from BR2 Q2 = 84oral 84